OpenStudy (anonymous):

Find the surface area of a square pyramid that has a base length of 14m and slant height of 19m

OpenStudy (anonymous):

please help

OpenStudy (mokeira):

\[A=a ^{2}+2a(\sqrt{\frac{ a ^{2} }{ 4 }+h ^{2}}\] \[=14^{2}+2(14)(\sqrt{\frac{ 14^{2} }{4 }+19^{2}}\] \[=196+28\times \sqrt{\frac{ 196 }{4 }+361}\] \[=196+28(\sqrt410)\] \[=196+28\sqrt{410}\] =762.96(to 2 decimal places)

OpenStudy (anonymous):

thank you @Mokeira

OpenStudy (mokeira):

u welcome :)